OpenStudy (anonymous):

Does anyone take Geometry A through Connections academy? I need help with a quiz

OpenStudy (dumbcow):

shoot whats the question

OpenStudy (juakikun):

Anyone found swapping answers for Connections Academy assignments (or any other school) will simply be suspended. Connections Academy does not even want their questions posted, so if you need help; try to post a similar or modified question to get help with. @amistre64 He posted this for a reason. Sorry but, We can't help you X:

OpenStudy (juakikun):

Plus thats the original question :c you should know the honor code from CA.

OpenStudy (anonymous):

so how would they like me to understand this....cause right now I dont, I just need it explained.

OpenStudy (juakikun):

If you could find a similar question in the math book, we can actually help on that and explain it o.o

OpenStudy (dumbcow):

here you go... the centroid of triangle is located 1/3 distance from midpoint and 2/3 from vertex  now apply this to your problem...you can still get plenty of help just dont copy/paste exact question from quiz

OpenStudy (dumbcow):

@Mel98

OpenStudy (anonymous):

Okay, but how do you figure it out

OpenStudy (dumbcow):

use the concept and apply to the specific problem you are given AC and BC in terms of "x" ... note AC is twice as long as BC you have to find length AB

OpenStudy (anonymous):

Yes

OpenStudy (dumbcow):

AC = 2BC  then you would say x-1 = 2(x-2) x-1 = 2x -4 3 = x total length of median AB = AC + BC AB = (x-1) + (x-2) AB = 2x -3 sub in x = 3 AB = 2(3) -3 AB = 6-3 = 3 now can you follow that example and solve your probelm